The consumer price index (CPI) rose from 105.9 to 113.4, while Helen's net salary decreased 7.0 %. Calculate the relative real change in Helen's salary (purchasing power) as percentage.

1
Expert's answer
2022-02-14T16:01:42-0500

113.4-105.9=7.5,

7.5/105.9=0.070821, 0.070821 *100%=7.0821%


7.0821%-7.0%=0.0821%